OpenStudy (anonymous):

f f(x) = 2x2 + 2, find f(5). 102 52 22 −1.5

OpenStudy (anonymous):

@jazy can you please help me!

OpenStudy (anonymous):

is this what u wrote up there?\[f(f(x))=2x^2+2\]

OpenStudy (anonymous):

yes @mukushla

OpenStudy (anonymous):

what math are u in?

OpenStudy (anonymous):

Algebra I

OpenStudy (anonymous):

maybe its\[f(x)=2x^2+2\]????

OpenStudy (anonymous):

\(2(5)^2+2 = 2 \times 5 \times 5 + 2\)

OpenStudy (anonymous):

Definitely @mukushla

OpenStudy (anonymous):

Not quite sure why you're struggling with this one @broweyebrit Do you not understand exponents, or functions or order of operations?

OpenStudy (anonymous):

yeah it is @mukushla